OpenStudy (anonymous):

Urgent Help Please A high school dance committee of 17 members is divided into the following subcommittees: Decorations: 4 members Music: 2 members Refreshments: 3 members Publicity: 5 members Ticket sales: 3 members. In how many different ways can the subcommittees be assigned?

OpenStudy (kropot72):

There are 17C4 ways of choosing the Decorations subcommittee. For each of these 17C4 ways there are 13C2 ways of choosing the Music subcommittee. Therefore there are \[17C4\times13C2=\frac{17!}{4! \times 13!} \times\frac{13!}{2!11!}\] ways of choosing the Decorations and Music subcommittees. Continuing this reasoning, the number of different ways of assigning the subcommittees is \[17C4\times13C2\times11C3\times8C5\times3C3=you\ can\ calculate\]

OpenStudy (anonymous):

Thank You =)

OpenStudy (kropot72):

You're welcome :)